OpenStudy (anonymous):

Quick review help for Trig? 2 questions?

OpenStudy (anonymous):

First one: What is the correct order of the functions from least to greatest according to the rate of change on the interval from x = -1 to x = 3? f(x), g(x), h(x) g(x), f(x), h(x) h(x), g(x), f(x) g(x), h(x), f(x) Here is the graph/charts: Also, I think the answer is C, but I'm really not sure, which is why I'm asking.

OpenStudy (anonymous):

Hold on, it didn't attach the whole picture.

OpenStudy (anonymous):

f(x) = (x + 3)^2 - 2, g(x) is the graph, and I will post h(x) in just a second.

OpenStudy (ipwnbunnies):

Ok, for g(x), the rate of change is constant because it is linear. The rate of change will be the slope. Just a little note.

OpenStudy (anonymous):

h(x) are the points: (-3,2) (-2,7) (-1,14) (0, 23) (1,34) (2,47) (3,62)

OpenStudy (ipwnbunnies):

Ok. Let's do the easy one first. Can you figure out the slope of g(x)?

OpenStudy (anonymous):

I believe it's 1/3, but I'm also thinking it's 3/1. I'm just really stressing out right now, I'm sorry.

OpenStudy (ipwnbunnies):

It's cool. Yes, the slope is 1/3 for g(x). Therefore, the rate of change og g(x) is 1/3. We'll make a note of that.

OpenStudy (ipwnbunnies):

Now, let's find the rate of change of h(x). We'll need a formula for that: \[avg. rate of change = \frac{f(b) - f(a)}{b-a}\]

OpenStudy (ipwnbunnies):

Doesn't this look like the slope formula? It's basically what it is. What is the slope of a line that joins too points on a graph. So, for h(x), what's the values of the function when x = -1 and 3?

OpenStudy (anonymous):

It does. Oh, so I need to find the slope between the points (-1, 14) and (3, 62), right?

OpenStudy (anonymous):

I got 12.

OpenStudy (ipwnbunnies):

Correct. That will be the average rate of change of the h(x) on that interval.

OpenStudy (ipwnbunnies):

Cool, I don't have a handy calculator, so I'll take your word lol. Now we have: rate of change of g(x) = 1/3 rate of change of h(x) = 12 Now we need the rate of change of f(x).

OpenStudy (ipwnbunnies):

There only gave us the function, though. We need to find the value at x = -1 and 3. Then, use the formula again to find the average rate of change.

OpenStudy (anonymous):

My iTouch is lost amongst all of my trig papers lol. I used the Google calculator.

OpenStudy (ipwnbunnies):

lol

OpenStudy (anonymous):

Hmm, I can't remember how I need to do that.

OpenStudy (anonymous):

Could I graph it and find the values that way?

OpenStudy (ipwnbunnies):

f(x) = (x+3)^2 - 2 Plug in -1 for x to find the value. Plug in 3 for x to find that value.

OpenStudy (ipwnbunnies):

No need for graphing.

OpenStudy (anonymous):

Oh, I'm sorry. For some reason I always take the hard route.

OpenStudy (anonymous):

For x = -1 I got the point (-1, 2). Let me check the second one real quick.

OpenStudy (ipwnbunnies):

That is correct.

OpenStudy (anonymous):

For x = 3, I got the point (3, 34). Now do I do the same thing I did before by subtracting the points, like in the formula you posted earlier?

OpenStudy (ipwnbunnies):

Yep :3

OpenStudy (anonymous):

Okay, I got 8.

OpenStudy (anonymous):

So the answer would be B, right?

OpenStudy (ipwnbunnies):

So did I in my head lol. This is what we have now: rate of change f(x) = 8 ROC g(x) = 1/3 ROC h(x) = 12

OpenStudy (ipwnbunnies):

Correct.

OpenStudy (anonymous):

Thank you so much for going through that with me. Would you have time to help me with one more?

OpenStudy (ipwnbunnies):

Sure.

OpenStudy (anonymous):

Thank you! :) For this question, I need to find which function has the most y-intercepts. I have two graphs, one for g(x) & h(x), and for f(x) I'll type it out real quick before I attach the graphs.

OpenStudy (ipwnbunnies):

A graph would be the most helpful for this lol.

OpenStudy (anonymous):

\[f(x) = x ^{3} + 2x ^{2} - 4x - 3\]

OpenStudy (anonymous):

g(x)

OpenStudy (anonymous):

h(x)

OpenStudy (anonymous):

For this one, I'm leaning towards f(x) as a guess.

OpenStudy (ipwnbunnies):

The most y-intercepts, or x-intercepts?

OpenStudy (anonymous):

Most x-intercepts. I'm sorry about that.

OpenStudy (ipwnbunnies):

Ok, let's look at the graphs first. An x-intercept is where the graph crosses the x axis. Let's look at g(x). How many times do you see it cross the x axis?

OpenStudy (anonymous):

It crosses 4 times.

OpenStudy (anonymous):

And for h(x), it crosses 3 times.

OpenStudy (ipwnbunnies):

Good, let's make a note for both of them. Now, f(x). \[f(x) = x ^{3} + 2x ^{2} - 4x - 3\]

OpenStudy (ipwnbunnies):

It would be best to graph it. If not, we would have to solve for x when y = 0. Remember, x-intercepts occur at y = 0.

OpenStudy (anonymous):

I have GeoGebra up, so I will be back in just a second.

OpenStudy (anonymous):

It also crosses the x-axis three times, right about the areas of x = -3, x = -0.75, and x = 1.5.

OpenStudy (anonymous):

So the answer would be g(x), correct?

OpenStudy (ipwnbunnies):

Correct.

OpenStudy (anonymous):

Thank you so much! I really appreciate you taking the time to help me!

OpenStudy (ipwnbunnies):

No problem.

OpenStudy (anonymous):

Have a good afternoon, and take care! :)